A 220v 60hz of alternating current is connected to an inductance of 3.5h what is the current passing through the inductor

1
Expert's answer
2022-03-09T15:54:44-0500

Solution.

"V=220V;"

"f=60Hz;"

"L=3.5H;"

"I-?;"

"I=\dfrac{V}{X_L};"

"X_L=2\pi f L;"

"X_L=2\sdot3.14\sdot60\sdot3.5=1318.8\Omega;"

"I=\dfrac{220}{1318.8}=0.167A;"

Answer: "I=0.167A."